OpenStudy (anonymous):

7e^x*sqrt(x) Thanks in advance for any help

OpenStudy (anonymous):

what do you want us to do with it? ^_^

OpenStudy (anonymous):

Do you want us to differentiate this equation?

OpenStudy (anonymous):

If so, Here is my solution. d/dx [ 7 e^x * sqrt(x) ] . . . apply the product rule 7 e^x * d/dx [ x^(1/2) ] + d/dx [ 7 e^x ] * x^(1/2) . . . d/dx x^(1/2) = 1/2 x^(-1/2) = 1 / (2 sqrt(x)) . . . d/dx 7 e^x = 7 e^x 7 e^x * 1 / (2 sqrt(x)) + 7 e^x * sqrt(x) 7 e^x / (2 sqrt(x)) + 7 e^x sqrt(x) . . . find a common denominator and combine 7 e^x / (2 sqrt(x)) + 7 e^x sqrt(x) 2 sqrt(x) / (2 sqrt(x)) (7 e^x + 7 e^x sqrt(x) 2 sqrt(x) ) / (2 sqrt(x)) (7 e^x + 7 e^x * 2x ) / (2 sqrt(x)) 7 e^x ( 1 + 2x) / (2 sqrt(x))

OpenStudy (anonymous):

Aha, thanks so much @Euler271 and @GirlByte. The command "differentiate" is so burned into my mind right now I forgot to include it. Thanks for doing this in steps, @GirlByte. This is a big help!